A strip of invisible tape 0.17 m long by 0.011 m w

Please answer in vector format! Thanks!!
Answer #1

a) Charge of the piece = 4/3 *10^-9 C = 1.333e-9 C

distance of the midpoint from origin = 0.17/3 m = 0.0567 m

Electric field at A = kq/r^2

= 9e9*1.333e-9/[0.0567^2 +0.09^2] *0.0567/sqrt[0.0567^2 +0.09^2] i + 9e9*1.333e-9/[0.0567^2 +0.09^2] *0.09/sqrt[0.0567^2 +0.09^2] j

= 565 i + 897j

Hence field is (565, 897, 0) N/C

b) Charge of middle piece = 4/3 *10^-9 C = 1.333e-9 C

distance of the midpoint from origin = 0 m

Electric field at A = kq/r^2

= 9e9*1.333e-9/[ 0.09^2]

=1481 j N/C

Hence field is (0, 1481, 0) N/C

part c) because of symmetrical to part a

E =  (-565, 897, 0) N/C

part d) Enet= [(565+0-565), (897+1481+897) , 0]

=  [ 0 , 3275 , 0 ] N/C